OpenStudy (anonymous):

Determine the points at which the following function is discontinuous and state the type of discontinuity: x^3-x/x(x^2-2x-3)

OpenStudy (anonymous):

Can you factor the denominator? That's where the discontinuities are. I'll display it in a moment and but I'd like it if you could try that yourself first and double-check to see if you didn't mean this instead: \[\frac{x^3-x}{x(x^2-2x-3)}\] What you wrote is the same as: \[{x^3-1*(x^2-2x-3)}\]

OpenStudy (anonymous):

It makes a huge difference in the answers so please check what it was you actually meant

OpenStudy (anonymous):

sorry, i should have mentioned when i put a "/" it means I'm dividing from that point :P i still don't know how to do the equations and everything on here.

OpenStudy (anonymous):

and i know 0/0 = asymptotic discontinuity and C/0 = hole discontinuity, its just a matter of actually finding them ;D haha